How much pure water must be mixed with 3 pints of 90% developer to produce a mixture that is 17% developer?
Found 2 solutions by Earlsdon, checkley77:
Answer by Earlsdon(6294)

Let x = the number of pints of water to be mixed with the 3 pints of 90% developer to obtain (3+x) pints of the 17% developer mixture.
3(0.9) = (3+x))(0.17) Simplify.
2.7 = 0.51+0.17x Subtract 0.51 from both sides.
2.19 = 0.17x Divide both sides by 0.17
12.8824 = x
Yuo need to mix 12.9 pints of water with the 3 pints of 90% developer mixture to obtain (12.9+3 = 15.9) pints of 17% developer mixture.

Answer by checkley77(12844)

.90*3=.17(3+x)
2.7=.51+.17x
-.17x=.51-2.7
-.17x=-2.15
x=-2.15/-.17
x=12.88 pints of water are added.
Proof:
.90*3=.17(3+12.88)
2.7=.17*15.88
2.7~2.7